Find the critical value za/2 Subscript alpha divided by 2 that corresponds to a = 0.06.

za/2 =
Answer #1

solution

the confidence level the z is ,

=0.06

/ 2 = 0.06 / 2 = 0.03

Z/2 = Z0.03 = 1.88 ( Using z table )